OpenStudy (anonymous):

Can someone please help me simplify this problem? (5x^3y)^2(-2x^5y^1)

OpenStudy (jhannybean):

\[(5x^{3y})^2\cdot (-2x^{5y})\] Is this it?

OpenStudy (jhannybean):

Or is it: \((5x^3y)^2\cdot (-2x^5y^1)\)

OpenStudy (anonymous):

It's the bottom one! :)

OpenStudy (anonymous):

Second one

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

First, distribute the square to both variables in your first function.\[(5)^2(x^3)^2(y)^2 \cdot (-2x^5y^1)\]

OpenStudy (jhannybean):

What do you get on the left side?

OpenStudy (anonymous):

25x^9y^2?

OpenStudy (anonymous):

\[25x ^{9}-y ^{2}?\]

OpenStudy (anonymous):

Sorry positive y

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

Almost. Try again. \((x^3)^2 = x^{3\cdot 2}\)

OpenStudy (anonymous):

x\[x^6?\]

OpenStudy (anonymous):

@Jhannybean

OpenStudy (anonymous):

\[25x ^{6}+y ^{2}\]

OpenStudy (anonymous):

?

OpenStudy (jhannybean):

Yes :) So you get. \(25x^6y^2 \cdot (-2x^5y^1)\) Now the next step.

OpenStudy (anonymous):

:-/ :(

OpenStudy (anonymous):

Oh umm think i got it!

OpenStudy (anonymous):

=D

OpenStudy (jhannybean):

Stay with me here. Now we have to group our like-terms.

OpenStudy (anonymous):

Would u multiply?

OpenStudy (anonymous):

=D

OpenStudy (jhannybean):

\[(25 \cdot (-2))(x^6\cdot x^5)(y^2\cdot y^1)\]\[=-50x^{6+5}y^{2+1}\] Simplify the powers and you're done!

OpenStudy (anonymous):

So It'd be -50x^11+ y^3 ?

OpenStudy (jhannybean):

Yes :)

OpenStudy (anonymous):

Omg!!!

OpenStudy (jhannybean):

Wait, no.

OpenStudy (jhannybean):

there is no addition involved here, all the variables are simply being multiplied.

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

:(

OpenStudy (jhannybean):

hollandlife Best Response Medals 0 So It'd be -50x^11+ y^3 ? 2 minutes ago There is no addition beween x^11 and y^3. It is simply -50x^11y^3

OpenStudy (jhannybean):

The only things that were added were the exponents, and that is because you had two like terms (i.e to x's) multiplying eachother with different powers.

OpenStudy (jhannybean):

Do you understand?